Find the area bounded by the curve y = |x−1|, the x-axis and the lines x = −7 and x = 11
1
Expert's answer
2020-06-08T20:58:49-0400

"\begin{array}{l}
A=\int_{a}^{b}|F(x)| d x 
=\int_{-7}^{11}|x-1| d x \\[1 em]
\text { It is clear from The drawing that } 
\text { integration must be divided into } 
\text { two parts } \\[1 em]
\begin{array}{l}
A=\int_{-7}^{1}(-x+1) d x+\int_{1}^{11}(x-1) d x \\[1 em]
=\left.\left(-\frac{x^{2}}{2}+x\right)\right|_{-7} ^{1}+\left.\left(\frac{x^{2}}{2}-x\right)\right|_{1} ^{11} \\[1 em]
=\left(-\frac{1}{2}+1-\left(-\frac{49}{2}-7\right)\right)+\left(\frac{121}{2}-11-\left(\frac{1}{2}-1\right)\right) \\[1 em]
=32+50 =82\\[1 em]
\text { You can find The area of two }
\text { tringles and add them } \\[1 em]
\begin{array}{l}
A=\frac{1}{2} \times 10 \times 10+\frac{1}{2} \times 8 \times 8 \\[1 em]
=50+32 \\[1 em]
=82
\end{array}
\end{array}
\end{array}"